# Catch Me if You Can: Detecting Unauthorized Data Use in Deep Learning Models

### v1

@@ 3.1.1 User capability
First, we assume $U$ can only mark a small number of samples, and these are to be included as part of a larger dataset. This simulates the realistic scenario where the dataset is curated from multiple data sources (e.g., [19]), and thus data contributed by each user constitutes only a small portion of the dataset (e.g., 0.1%).
Next, in the provenance verification process, we assume $U$ has access to an auxiliary set of non-member samples, which is a common assumption to control the MI process at a low FPR regime [8, 6, 70]. $U$ also has black-box access to query the target model. In a real world setting, users often have limited control and knowledge beyond their own data, and thus we specify the constraints by $U$ as follows.
  • $U$ cannot modify their data’s labels, which are assigned by the model owner or an external service [21, 68, 19].
@@
  • $U$ has no expertise/resources to train any shadow/proxy models, as in many settings, training even a single shadow model can impose prohibitively high data and compute requirements (especially for the large models) [65].
3.1.2 Evaluation metric
We follow the best practice to measure the MI success by its true positive rate (TPR) and false positive rate (FPR) [8]. There are two ways to define the success metric: 1g11g TPR under a fixed FPR; or 1g21g FPR under a fixed TPR. We use metric 1g21g, and we first explain why 1g11g is unsuitable for evaluating the data-provenance based MI method in our work.
Motivation. Metric 1g11g is commonly adopted by existing work where MI is posed as an attack to expose privacy leakage in DL models [8, 6, 70]. Their goal is to evaluate whether the adversary can reliably de-identify (even just a few of) the training samples. This places emphasis on the requirement of low FPR, for which 1g11g represents a suitable metric [8, 6, 70].
In comparison, in our work of tracing data provenance, both low FPR (to avoid false accusation) and high TPR (for correct data provenance) are crucial. However, manually choosing a low FPR threshold and evaluating the corresponding TPR (as in 1g11g) can be undesirable, as the pre-defined FPR threshold may be too conservative or aggressive. We use Fig. 2 (from our experiment) to explain next.
For the left method in Fig. 2, using a 0.1% FPR threshold is too conservative: this method can yield the same 100% TPR under a lower FPR of 0%. Setting a 0% FPR threshold is appropriate for the left method, but this becomes too aggressive for the right method in Fig. 2: this other high-power method (with 0.9993 AUC) yields a poor outcome of 5% TPR.
Figure 2: Comparing two metrics for evaluating MI success in our work: 1g11g TPR@fixed FPR; 1g21g FPR@fixed TPR.
@@ 3.1.2 Evaluation metric
Instead, MI for data provenance requires both low FPR and high TPR, for which 1g21g is a more suitable evaluation metric.
@@
Under a 100% TPR, the left method in Fig. 2 has a 0% FPR, and the right method has a slightly higher 0.52% FPR. Both results are able to appropriately reflect the two MI methods’ capability in fulfilling the requirement of low FPR (0% and 0.52%) [^1] Note that this metric serves only to evaluate the MI success, and the users can still choose an arbitrary FPR threshold (e.g., 0% or 1% FPR) to control the MI process when auditing their actual data (Section 4.4.3). and high TPR (100%) for tracing data provenance.
Based on the above, we adopt FPR@fixed TPR as the evaluation metric in this work, and we set a 100% TPR threshold to simulate the case where all the users whose data are misused without their consent can trace the provenance of their data (see our design goal in Section 4 next). An ideal technique should incur as low FPR as possible.
4 Methodology
We first describe our design goals, and then conduct an empirical study to understand why directly leveraging existing MI methods cannot meet our goals. We then present MembershipTracker, a two-step technique with a data marking and provenance verification process: first by marking the users’ protected data with targeted changes (Section 4.3); then performing a specialized MI process to audit whether the marked data are used to train the model (Section 4.4).
@@ 4 Methodology
4.1 Motivation
Our work addresses data provenance through MI, and there are today a number of available MI methods [70, 6, 8, 65]. However, these methods still suffer from limited MI effectiveness and/or impose strong assumptions on the users - we elaborate both limitations next.
@@ 4.1 Motivation
To validate this, we use two methods in the state-of-the-art Likelihood-ratio attack, one is with and the other without shadow model calibration (128 shadow models) [8]. We evaluate both methods on a WideResNet-28-4 model trained on half of the training set in CIFAR100 dataset, and they both incur a high FPR of 75.32% and 64.62%, under 100% TPR.
With that said, the above results should not be interpreted as the failure of these approaches, because they were originally built from an adversarial perspective to expose the worst-case privacy leakage in DL models, where even just reliably de-identifying a few member samples is considered meaningful [8], e.g., the shadow-model based method in Fig. 3 can achieve 27.21% TPR@0.1% FPR.
@@
As in Fig. 3, existing MI methods still struggle with this task. The reason is that the training samples are not always strongly memorized by the model (except some outliers) [71, 8, 65], which is directly related to the ability of performing accurate MI. This problem motivates related work [65, 13, 10] to amplify the model’s memorization on the training samples, and we discuss their limitation next.
Figure 3: MI for tracing data provenance requires both high TPR (for correct provenance detection), and low FPR (to avoid false accusation). However, even state-of-the-art MI methods [8] cannot meet this challenging goal, and are unsuitable for data provenance purposes.
@@
State-of-the-art MI methods either require training shadow models or assume control to manipulate the data labels, both are beyond ordinary users’ capabilities.
For example, leading MI methods [8, 6, 70] commonly require training shadow models to calibrate the inference process, which assumes additional data and compute access, and is challenging for the non-expert users.
There are other methods that do not require shadow-model calibration, but they either suffer from poor effectiveness [71, 8] or impose unrealistic assumption on the users [65, 10]. For instance, the approach by Tramer et al. [65] can improve the MI success without shadow-model calibration, but it instead assumes the users have the ability to mislabel some training data, which can be challenging in the cases where the users have no control over the data labels [21, 68, 19]. Even if this is feasible, we find that the increase of MI success is still largely limited (see Appendix E).
4.2 Overview
@@
4.3 Data Marking
The goal of data marking is to ensure that the membership of the marked samples can be reliably de-identified. To this end, we first observe that the ability to perform accurate MI is directly related to the model’s propensity in memorizing individual data points, and data that are strongly memorized by the model (e.g., outlier samples, mislabeled samples) are known to be easier to be de-identified [8, 71, 65, 14]. Based on this, our goal can be formulated as: how can the users modify their data to amplify the model’s memorization on the target data while preserving high visual quality?
@@ 4.3 Data Marking
4.3.1 An initial strawman approach
Inspired by the observation that atypical samples presented in the dataset are easier to be de-identified (e.g., mislabeled samples [65], or samples that are out-of-distribution - OOD [8]), a straightforward approach is to directly replace the original target samples as OOD samples.
There are different ways to generate OOD samples, and we develop a simple method that creates samples consisting of random color stripes, such as the one shown below (the method’s details are deferred to Section 5). Meanwhile, other alternative methods such as using samples from an OOD dataset may also be considered (Appendix C.1.3).
We first conduct an experiment to validate the effectiveness of the above approach. We use the CIFAR100 dataset with half of the training set for model training. We consider 100 target users from different classes replacing their data as some random OOD samples, and each contributes a small number of samples to the training set (0.1% in proportion). In terms of the signal function for MI, we follow prior work [54, 70] to use the prediction loss values.
@@ 4.3.1 An initial strawman approach
4.3.2 Step 1: Image blending
As discussed earlier, the marked samples should contain the original feature (to preserve visual information), as well as the OOD feature (for provenance purposes). Inspired by the common image blending technique [74, 68], we propose to blend the target sample with the OOD feature:
$x\oplus(x_{ood},m)=m\cdot x+(1-m)\cdot x_{ood},$ (1)
@@ 4.3.2 Step 1: Image blending
Figure 4: The two-step data marking process: (1) blend the original samples with OOD feature; (2) inject procedural noise. These subtle changes can induce the target samples to be strongly memorized by the model trained on them.
4.3.3 Step 2: Noise injection
Our idea is to inject a small amount of noise to suppress the influence of the original feature, and have the model memorize the OOD feature.
For this, we draw inspiration from the concept of adversarial samples, which works by adding imperceptible noise to the inputs and cause them to be misclassified by the model [43, 40, 17]. The perturbation can thus be viewed as the noise that can suppress the influence of the original feature and cause the model to predict the wrong label.
In our context, we can inject adversarial perturbation into the users’ target samples and prompt the model to memorize the OOD feature. There are different ways to generate the perturbation. Optimization-based approach requires access to some proxy model to generate the perturbation [43, 40], which can still be challenging for the ordinary users we consider.
We therefore resort to another optimization-free strategy that does not assume any additional access. In particular, Co et al. [17] propose that procedural noise functions can be used to generate input-agnostic adversarial perturbation. These functions are commonly used in computer graphics to generate different textures and patterns and enrich the image details [17]. Co et al. find that the procedural noise patterns share visual resemblance with existing adversarial perturbation patterns, and demonstrate that they can be similarly used to construct adversarial samples.
This inspires our solution in injecting perlin noise [17] to the target samples, and its generation is as follows.
$S_{perlin}(x,y)=\sum_{n=1}^{\Omega}p(x\cdot\frac{2^{n-1}}{\lambda_{x}},y\cdot\frac{2^{n-1}}{\lambda_{y}}),$ (2)
$G_{perlin}(x,y)=sin((S_{perlin}(x,y))\cdot 2\pi\phi_{sine}),$ (3)
where $\lambda_{x},\lambda_{y},\Omega,\phi_{sine}$ are for controlling the noise value at point $(x,y)$. $\lambda_{x},\lambda_{y}$ are the wavelengths, $\Omega$ is the number of octaves that contribute to the most visual change, $\phi_{sine}$ is the periodicity of the sine function that creates distinct bands in the image to achieve a high frequency of edges, which can achieve more distinct visual patterns [17]. We follow Co et al. to use the $L_{\infty}$ norm as the perturbation budget $\delta$ and create random perlin noise [17] for data marking.
@@ 4.3.3 Step 2: Noise injection
Overall, the two-step marking process is able to reduce the FPR from 56.32% to 7.97%. While the FPR can be further reduced by using a smaller $m$ for image blending or larger $\delta$ for noise injection, it would undesirably increase the visual distortion (see Appendix C.1.2 for an evaluation). We thus propose another innovation in the MI process to mitigate the FPR, and without degrading the image quality.
4.4 Data Provenance via Membership Inference
To start with, we note that so far we have been considering MI on a common instance basis [56, 8, 6, 70], which uses the per-instance loss as the signal function value for MI. We first explain why this approach would incur high FPR, and then present a specialized MI process to overcome it.
4.4.1 Why instance-based MI suffers from high FPR
@@
For instance, the left figure in Fig. 5 shows the individual loss values on different samples, where many non-member samples also yield low prediction loss (the low loss values are shown towards the right side because they are plotted in the logit-scaled form [8] for visualization purposes). This consequently leads to a 7.97% FPR @100% TPR.
Figure 5: Comparison of using the per-instance loss and the average loss across the samples by each user (each contributes 0.1% of the training data) as the signal function for MI verification. The former yields 7.97% FPR@100% TPR; while the latter has 0% FPR@100% TPR (our proposal).
4.4.2 Set-based membership inference for reliable data provenance
To overcome the above, our key insight is that, even though each target user only contributes a small number of samples to the model’s training set (e.g., 25 out of 25,000 instances), they can still leverage the collective information across their own samples to construct a more reliable signal function for MI, thereby reducing the FPR.

### v2

@@ 3.1.1. User capability
First, we assume $U$ can only mark a small number of samples, and these are to be included as part of a larger dataset. This simulates the realistic scenario where the dataset is curated from multiple data sources (e.g., (Deng et al., 2009)), and thus data contributed by each user constitutes only a small portion of the dataset (e.g., 0.1%).
Next, in the provenance verification process, we assume $U$ has access to a set of non-member samples that are not used for training the target model [^1] Under the presumption that only a small fraction of data sampled from the distribution were used in training, one may simply draw a random sample from the underlying distribution, and be confident that it is representative of the non-member data (Bertran et al., 2024; Ye et al., 2022)., which is a common assumption [^2] While we also adopt this common assumption in our study, we conduct additional experiments and find that this assumption can in fact, be relaxed (see Section 7). to control the MI process at a low FPR regime (Carlini et al., 2022; Bertran et al., 2024; Ye et al., 2022). $U$ also has black-box access to query the target model. In a real world setting, users often have limited control and knowledge beyond their own data, and thus we specify the constraints by $U$ as follows.
  • $U$ cannot modify their data’s labels, which are assigned by the model owner or an external service (Dzieza, 2023; Wenger et al., 2024; Deng et al., 2009).
@@
  • $U$ has no expertise/resources to train any shadow/proxy models, as in many settings, training even a single shadow model can impose prohibitively high data and compute requirements (especially for the large models) (Tramèr et al., 2022).
3.1.2. Evaluation metric
We follow the best practice to measure the MI success by its true positive rate (TPR) and false positive rate (FPR) (Carlini et al., 2022), which is fully characterized by the Receiver Operating Characteristic (ROC) curve. From the ROC curve, we can compute either 1g11g TPR under a fixed FPR; or 1g21g FPR under a fixed TPR. We use metric 1g21g, and we first explain why 1g11g is unsuitable for evaluating the data-provenance based MI method in our work.
Motivation. The goal of performance evaluation is to understand whether MembershipTracker can achieve both low FPR (to avoid false accusation) and high TPR (for correct data provenance). However, manually choosing a low FPR threshold and evaluating the corresponding TPR (as in 1g11g) can be undesirable, as the pre-defined FPR threshold may be too conservative or aggressive. We use Fig. 2 (from our experiment) to explain the tradeoffs involved.
For the left method in Fig. 2, using a 1% FPR threshold is too conservative. Under this threshold, both methods in Fig. 2 will be deemed to have the same performance, yet the left method is in fact much more powerful and it can yield the same 100% TPR under a lower FPR of 0% (undesirable).
Likewise, setting a tighter FPR threshold such as 0%, is appropriate for the left method, but this becomes too aggressive for the right method in Fig. 2 - this high-power method (with 0.9993 AUC) yields a poor outcome of 5% TPR (undesirable).
Figure 2. From the same ROC curve, we can compute either 1g11g TPR@fixed FPR; or 1g21g FPR@fixed TPR.
@@ 3.1.2. Evaluation metric
Instead, metric 1g21g is able to properly reflect the high performance by both methods (with low FPRs and high TPRs), and thus we adopt it for performance evaluation purposes.
@@
Under a 100% TPR, the left method in Fig. 2 has a 0% FPR, and the right method has a slightly higher 0.52% FPR. This result shows that (1) both methods have very good performance (in having low FPRs and high TPRs); and (2) the left method is slightly better than the right one (0% vs 0.52% FPR).
Based on the above, we adopt FPR@fixed TPR as the evaluation metric, and we set a 100% TPR threshold to simulate the case where all the users whose data are misused without their consent can trace the provenance of their data (see our design goal in Section 4). An ideal technique should incur as low FPR as possible.
Remark. Note that this metric serves to measure the minimal FPR for MembershipTracker to achieve 100% TPR. We will show later that MembershipTracker has very low FPR@100% TPR. Therefore, this result represents a desirable FPR threshold that can be used to calibrate the actual MI data auditing process, which follows the common practice to control the MI process under the low FPR regime, and detects whether the testing data was used for training or not (details in Section 4.4.3).
For instance, if MembershipTracker achieves 0.1% FPR@100% TPR, it means users can control the MI process under 0.1% FPR [^3] Note that users can still choose an arbitrary FPR threshold (e.g., 0% or 1% FPR)., and MembershipTracker will be able to detect those data used for training (with 100% TPR).
In essence, metric 1g11g and metric 1g21g merely represent specific points in the ROC curve, and a full performance characterization should still be captured by the ROC curve. However, for MembershipTracker, we find that using metric 1g21g is already able to illustrate that MembershipTracker can achieve high performance in the low FPR regime. For example, our evaluation in Section 5.1 shows that MembershipTracker achieves an average of 0% FPR@100% TPR across different evaluation settings, which means MembershipTracker can achieve perfect discrimination in the low FPR regime. In these cases, all ROC curves will show a flat line similar to that on the left of Fig. 2, and thus we omit the curves for space reasons.
With that said, in evaluating other related techniques (Carlini et al., 2022; Wenger et al., 2024; Tramèr et al., 2022), we have carefully provided the full ROC curve for a comprehensive comparison (e.g., Fig. 3, Fig. 15, and Fig. 17).
4. Methodology
We first describe our design goals, and then conduct an empirical study to understand why directly leveraging existing MI methods cannot meet our goals. We then present MembershipTracker with a data marking and provenance verification process: first by marking the users’ protected data with targeted changes (Section 4.3); then performing a specialized MI process to audit whether the marked data are used to train the model (Section 4.4).
@@ 4. Methodology
4.1. Motivation
Our work addresses data provenance through MI, and there are several available MI methods (Ye et al., 2022; Bertran et al., 2024; Carlini et al., 2022; Tramèr et al., 2022). However, these methods still suffer from limited MI effectiveness and/or impose strong assumptions on the users - we elaborate both limitations next.
@@ 4.1. Motivation
To validate this, we use two methods in the state-of-the-art Likelihood-ratio attack, one is with and the other without shadow model calibration (128 shadow models) (Carlini et al., 2022). We evaluate both methods on a WideResNet-28-4 model trained on half of the training set in CIFAR100 dataset, and they both incur a high FPR of 75.32% and 64.62%, under 100% TPR.
With that said, the above results should not be interpreted as the failure of these approaches, because they were originally built from an adversarial perspective to expose the worst-case privacy leakage in DL models, where even just reliably de-identifying a few member samples is considered meaningful (Carlini et al., 2022), e.g., the shadow-model based method in Fig. 3 can achieve 27.21% TPR@0.1% FPR.
@@
As in Fig. 3, existing MI methods struggle with this task. The reason is that the training samples are not always strongly memorized by the model (except some outliers) (Yeom et al., 2018; Carlini et al., 2022; Tramèr et al., 2022), which is directly related to the ability of performing accurate MI. This problem motivates related work (Tramèr et al., 2022; Chen et al., 2022; Chaudhari et al., 2023) to amplify the model’s memorization on the training samples, and we discuss their limitation next.
Figure 3. MI for tracing data provenance requires both high TPR (for correct provenance detection), and low FPR (to avoid false accusation). However, even state-of-the-art MI methods (Carlini et al., 2022) cannot meet this challenging goal, and are unsuitable for data provenance purposes.
@@
State-of-the-art MI methods either require training shadow models or assume control to manipulate the data labels, both are beyond ordinary users’ capabilities. For example, leading MI methods (Carlini et al., 2022; Bertran et al., 2024; Ye et al., 2022) commonly require training shadow models to calibrate the inference process, which assumes additional data and compute access, and is challenging for the non-expert users.
There are other methods that do not require shadow-model calibration, but they either suffer from poor effectiveness (Yeom et al., 2018; Carlini et al., 2022) or impose unrealistic assumption on the users (Tramèr et al., 2022; Chaudhari et al., 2023). E.g., the approach by Tramer et al. (Tramèr et al., 2022) can improve the MI success without shadow-model calibration, but it assumes users have the ability to mislabel some training data, which can be challenging in the cases where the users have no control over the data labels (Dzieza, 2023; Wenger et al., 2024; Deng et al., 2009). Even if this is feasible, we find that the increase of MI success is still largely limited (see Appendix E for a validation).
4.2. Overview
@@
4.3. Data Marking
The goal of data marking is to ensure that the membership of the marked samples can be reliably de-identified. To this end, we first observe that the ability to perform accurate MI is directly related to the model’s propensity in memorizing individual data points, and data that are strongly memorized by the model (e.g., outlier samples, mislabeled samples) are known to be easier to be de-identified (Carlini et al., 2022; Yeom et al., 2018; Tramèr et al., 2022; Chen and Pattabiraman, 2024a). Based on this, our goal can be formulated as: how can the users modify their data to amplify the model’s memorization on the target data while preserving high visual quality?
@@ 4.3. Data Marking
4.3.1. An initial strawman approach
Inspired by the observation that atypical samples presented in the dataset are easier to be de-identified (e.g., mislabeled samples (Tramèr et al., 2022), or samples that are out-of-distribution - OOD (Carlini et al., 2022)), a straightforward approach is to directly replace the original target samples as OOD samples.
There are various ways to generate OOD samples, and we develop a simple method that creates samples with random color stripes, such as the one shown below (the method’s details are in Section 5). Meanwhile, other alternative methods such as using samples from an OOD dataset may also be considered (Appendix C.1.3).
We first conduct an experiment to validate the effectiveness of the above approach. We use the CIFAR100 dataset with half of the training set for model training. We consider 100 target users from different classes replacing their data as some random OOD samples, and each contributes a small number of samples to the training set (0.1% in proportion). In terms of the signal function for MI, we follow prior work (Sablayrolles et al., 2019; Ye et al., 2022) to use prediction loss values.
@@ 4.3.1. An initial strawman approach
4.3.2. Step 1: Image blending
As discussed earlier, the marked samples should contain the original feature (to preserve visual information), as well as the OOD feature (for provenance purposes). Inspired by the common image blending technique (Zhang et al., 2018; Wenger et al., 2024), we propose to blend the target sample with the OOD feature:
(1) $x\oplus(x_{ood},m)=m\cdot x+(1-m)\cdot x_{ood},$
@@ 4.3.2. Step 1: Image blending
Figure 4. The two-step data marking process: (1) blend the original samples with OOD feature; (2) inject procedural noise. These subtle changes can induce the target samples to be strongly memorized by the model trained on them.
4.3.3. Step 2: Noise injection
Our idea is to inject a small amount of noise to suppress the influence of the original feature, and have the model memorize the OOD feature. For this, we draw inspiration from the concept of adversarial samples, which works by adding imperceptible noise to the inputs and cause them to be misclassified by the model (Liu et al., 2016; Li et al., 2020; Co et al., 2019). The perturbation can thus be viewed as the noise that can suppress the influence of the original feature and cause the model to predict the wrong label.
In our context, we can inject adversarial perturbation into the users’ target samples and prompt the model to memorize the OOD feature. There are different ways to generate the perturbation. Optimization-based approach requires access to some proxy model to generate the perturbation (Liu et al., 2016; Li et al., 2020), which can still be challenging for the ordinary users we consider.
We therefore resort to another optimization-free strategy that does not assume any additional access. In particular, Co et al. (Co et al., 2019) propose that procedural noise functions can be used to generate input-agnostic adversarial perturbation. These functions are commonly used in computer graphics to generate different textures and patterns and enrich the image details (Co et al., 2019). Co et al. find that the procedural noise patterns share visual resemblance with existing adversarial perturbation patterns, and demonstrate that they can be similarly used to construct adversarial samples.
This inspires our solution in injecting perlin noise (Co et al., 2019) to the target samples, and its generation is as follows.
(2) $S_{perlin}(x,y)=\sum_{n=1}^{\Omega}p(x\cdot\frac{2^{n-1}}{\lambda_{x}},y\cdot\frac{2^{n-1}}{\lambda_{y}}),$
(3) $G_{perlin}(x,y)=sin((S_{perlin}(x,y))\cdot 2\pi\phi_{sine}),$
where $\lambda_{x},\lambda_{y},\Omega,\phi_{sine}$ are to control the noise value at point $(x,y)$. $\lambda_{x},\lambda_{y}$ are the wavelengths, $\Omega$ is the number of octaves that contribute to the most visual change, $\phi_{sine}$ is the periodicity of the sine function that creates distinct bands in the image to achieve a high frequency of edges, which can achieve more distinct visual patterns (Co et al., 2019). We follow Co et al. to use $L_{\infty}$ norm as the perturbation budget $\delta$ and create random perlin noise (Co et al., 2019) for data marking.
@@ 4.3.3. Step 2: Noise injection
Overall, the two-step marking process is able to reduce the FPR from 56.32% to 7.97%. While the FPR can be further reduced by using a smaller $m$ for image blending or larger $\delta$ for noise injection, it would increase the visual distortion (see Appendix C.1.2 for an evaluation). We thus propose another innovation in the MI process to mitigate the FPR, and without degrading the image quality.
4.4. Data Provenance via Membership Inference
To start with, we note that so far we have been considering MI on a common instance basis (Shokri et al., 2017; Carlini et al., 2022; Bertran et al., 2024; Ye et al., 2022), which uses the per-instance loss as the signal function value for MI. We first explain why this approach would incur high FPR, and then present a specialized MI process to overcome it.
4.4.1. Why instance-based MI suffers from high FPR
@@
For instance, the left figure in Fig. 5 shows the individual loss values on different samples, where many non-member samples also yield low prediction loss (the low loss values are shown towards the right side because they are plotted in the logit-scaled form (Carlini et al., 2022) for visualization purposes). This leads to a 7.97% FPR @100% TPR.
Figure 5. Comparison of using the per-instance loss and the average loss across the samples by each user (each contributes 0.1% of the training data) as the signal function for MI verification. The former yields 7.97% FPR@100% TPR; while the latter has 0% FPR@100% TPR (our proposal).
4.4.2. Set-based membership inference for reliable data provenance
To overcome the above, our key insight is that, although each target user only contributes a small number of samples to the model’s training set (e.g., 25 out of 25,000 instances), they can still leverage the collective information across their own samples to construct a more reliable signal function for MI, and reduce FPR.